One endpoint of a segment is (12, -8). The midpoint is (3, 18). Find the coordinates of the other endpoint.

Answer by edjones(7311)

(12,-8)-(3,18)=(9,-26)
(3,18)-(9,-26)=(-6,44) the other end point.
Ed